OpenStudy (anonymous):

really need help. dont understand Given the points (2, 3) and (6, 8), write the point-slope form of the equation. A. y - 3 = x - 2 B. y - 3 = 5/4 (x - 2) C. y - 8 = 4/5 (x - 6)

OpenStudy (anonymous):

Do you know what point slope form is?

OpenStudy (anonymous):

Start by finding the slope between the two points: (2, 3) and (6, 8)

OpenStudy (anonymous):

Once you have the slope between the two points, use of the points and the point slope formula to create the proper equatioo.

OpenStudy (anonymous):

*equation

OpenStudy (anonymous):

how do i find the slope?

OpenStudy (anonymous):

Subbtract the y-values and then subtract the x values. Then divide the resultant y values over the x values.

OpenStudy (anonymous):

(2, 3) and (6, 8), the y values are 8 and 3. the x values are 6 and 2. So 8-3=5 and 6-2=4. Now divide 5/4... This is your slope.

OpenStudy (anonymous):

the slope is 5/4 or do i divide that first then i will have the slope?

OpenStudy (anonymous):

5/4 is the slope

OpenStudy (anonymous):

ok now what do i do next?

OpenStudy (anonymous):

now use the formula y-y1=m(x-x1). Replace m with the slope that you just found. and x1,y1 with one of teh coordinate points that were give.

OpenStudy (anonymous):

The annual interest rate of Jacques' savings account is 3.6%, and simple interest is calculated monthly. What is the periodic interest rate of Jacques' account?

OpenStudy (anonymous):

where would i put the coordinate that i choose?

OpenStudy (anonymous):

y-y1=m(x-x1). So replace m with 5/4.. This will look like y-y1=5/4(x-x1). Now say you chose to use the point (2, 3). This will look like y-3=5/4(x-2) If you were to use the point (6, 8), you will get something like y-8=5/4(x-6) both y-8=5/4(x-6) and y-3=5/4(x-2) are the same answers. It just will take some math to work them around to notice that.

OpenStudy (anonymous):

so it could be either one of those answers?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you so much for your time and all of your help. i truly appreciate it

OpenStudy (anonymous):

But look what choices you have in the multiple choice. The only one that matches that is B: y-3=5/4(x-2)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Yep